| Operation Summary for January 27, 2012 - January 30, 2012 |
|
NuMI reported receiving 8.93E18 protons on target for the period from 1/23/12 to 1/30/12. Friday January 27 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE and NuMI experiments taking beam. During the day shift, Operations turned off the Booster beam at 9:37 AM in preparation for an access to repair a RF station (BRF17). Operators loaded an energy conservation timeline into the Time Line Generator (TLG) at 9:41 AM. Operations reported at 10:41 AM that the MI gap-clearing kickers had been repaired. At 11:38 AM, Operators began preparing NuMI for access. Personnel at the Neutron Therapy Facility reported at 11 AM that they had treated their last patient for the day. The Booster RF work ended and Operations established beam to the Meson FTBF experiment T-1018 at 12:30 PM. Operations reestablished beam to MiniBooNE at 12:34 PM. NuMI came out of access and resumed taking beam at 1:24 PM. At 2:10 PM, the NuMI beam permit tripped off due to bad beam positions; it reset. At 2:28 PM a Linac Klystron RF station (KRF3) tripped off; it reset. Operations reported problems with the Linac output at 3:41 PM; Linac personnel investigated and found that the 750 KeV line emittance probe had ran itself into the beam. During the evening shift, at 5:13 PM, BRF13 tripped off; Operators had to reset it locally. A MI person completed a study at 5:27 PM that he had started on the day shift. At 5:36 PM, KRF1 tripped off; it reset. Operations disabled the T-1018 beam at 9:56 PM. Saturday January 28 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, KRF1 tripped off at 12:34 AM; it reset. At 1:43 AM, BRF4 tripped off for the first of four times in two and half hours; Operators lowered the anode program voltage for BRF4 after its last trip. Operators reported at 1:46 AM that the I- Source had spiked several times and caused NuMI trips. At 2:18 AM, BRF3 tripped off; it reset. At 4:29 AM, MIRF18 tripped off; it reset. During the day shift, Operations established beam to T-1018 at 10:27 AM. Operations reported at 11:20 AM that the output of the I- Source had been dropping. At 11:28 AM, the MI RF South Anode Power Supply tripped off; it reset. At 11:58 AM, a Booster vacuum pump (B:IPRF15) tripped off and took down BRF3 and 4 with it; everything recovered and reset. At 3:16 PM, a Klystron Debuncher tripped off and would not reset; Operators called in Linac personnel. During the evening shift, Operations prepared T-1018 for access at 8:07 PM. Operators loaded an energy conservation timeline into the TLG at 5:43 PM for the Linac repairs. The MI South Anode Power Supply tripped off again; it reset. Linac resumed sending beam at 9:39 PM. BRF5 tripped off at 9:39 PM; it reset. Operators completed a Booster injection tune at 11:58 PM. Sunday January 29 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 4:04 AM, three Booster RF stations (BRF14, 15, and 16) all tripped off; they all reset. At 5:31 AM, BRF4 tripped off for the first of two times in about 75 minutes; it reset. During the day shift, at 8:33 AM, BRF4 tripped off; it reset. Operations reported at 9:57 AM that I- Source spikes continued to cause NuMI beam permit trips; Operators made adjustment to Booster to limit the occurrences of the NuMI trips. Operations established beam to T-1018 at 10:41 AM. BRF15 tripped off at 12:15 PM; it reset. During the evening shift, BRF4 tripped off at 4:35 PM; it reset. Operator reported a problem with BRF1; a power cycle of its power supply resolved the problem. Operators reported a problem with BRF16; it reset. Operations disabled T-1018 beam at 9:59 PM. Monday January 30 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 12:32 AM, the I- Source’s high voltage tripped off; it reset. Operators reported getting repeated Enclosure C sump alarms; Operators accessed the enclosure at 7:08 AM and found a sump discharge line with a loose coupling that was spraying water back into the sump pit. The Duty Mechanic accessed Enclosure C at 7:48 AM to make repairs. Operations Coordinator (OpsCo) The OpsCo agrees with the plans by Muon Rings personnel to take low intensity beam for studies. A seven-day shutdown begins tomorrow, 1/31/12, to replace a MI quadrupole magnet. Beam will go off at 2 AM. MI keys will be available by 6 AM. Other machines and experiments will conduct maintenance that fits within the seven-day time frame. Linac will also conduct maintenance, but they will turn on to send beam to NTF when needed. |
|
More Information Comments and Suggestions |
| last modified 01/30/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |